# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ Abstract.
The aim of this article is twofold. First, we develop a unified framework for viscosity solutions to both first-order Hamilton-Jacobi equations and semilinear Hamilton–Jacobi equations driven by the idiosyncratic operator. Second, we establish a vanishing-viscosity limit—extending beyond the classical control-theoretic setting—for solutions of semilinear Hamilton–Jacobi equations, proving their convergence to the corresponding first-order solution as the idiosyncratic noise vanishes. Our approach provides an optimal convergence rate.
@@ 1. Introduction
(1.2) $\displaystyle-\partial_{t}U(t,\mu)\;+\;\int_{\mathbb{R}^{d}}H\bigl(x,\partial_{\mu}U(t,\mu,x),\mu\bigr)\,d\mu(x)\;-\;\int_{\mathbb{R}^{d}}\mathrm{tr}\!\left[a(\mu,x)\,\nabla_{x}\partial_{\mu}U(t,\mu,x)\right]d\mu(x)=0,$ $\text{ in }[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$
where, $T>0$, $H:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is the Hamiltonian, and $a:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to M_{d\times d}(\mathbb{R})$ is the diffusion matrix that will be assumed to be positive definite. Usually we will consider the final time condition $U(T,\cdot)=\mathcal{G}(\cdot)$ where $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is the terminal datum. Here, $\partial_{\mu}U(t,\mu,x)$ stands for the Wasserstein gradient of $U$ at the point $(t,\mu,x)$.
@@
In this paper, we develop an approach that enables us to treat these two equations using a unified definition of a viscosity solution. In particular, we establish an asymptotic relation between the viscosity solutions of (1.1) and (1.2) through a vanishing-viscosity limit. To the best of our knowledge, this is the first work to provide such a perturbative link without imposing regularity assumptions on the functions $U^{\varepsilon}$, the viscosity solutions to (1.2) when $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$, apart from those inherent to the control setting. In [tangpi2022probabilisticapproachvanishingviscosity], a vanishing viscosity result is obtained also in the presence of a common noise operator, in the convex case; however, that analysis relies on additional regularity assumptions on the solutions.
@@ 1.1. Overview on the existing literature
The study of HJEs on Wasserstein spaces has gained a lot of interest, in these last years, thanks to their connections with Mean Field Control, Mean Field Games, and the associated Master Equation. Mean Field Games theory has been introduced by Lasry and P.L.Lions [LLMeanFG], and independently by Huang, Caines and Malhamé [MeanFieldCaines]. It has been subsequently developed in Lions’ courses at Collège de France [College] and it has attracted the interest of a large group of researchers. The well-posedness theory for solutions of HJEs in the space of probability measures is one of the objectives of this theory due to its connection with mean-field particle approximation [cardaliaguet2022regularityvaluefunctionquantitative, cecchin2025quantitativeconvergencemeanfield], the control of Mckean-Vlasov equation [CarmonaDelarue, CarmonaDelarueLachapelle2013], and large deviation theory [FengKurtz2006].
In the study of HJEs in $\mathcal{P}_{2}(\mathbb{R}^{d})$, various notions of viscosity solutions have been proposed in the literature (possibly in presence of a common noise) with the shared aim of extending the classical finite-dimensional framework to this more complex setting. However, this new context presents significant challenges. The geometry of $\mathcal{P}_{2}(\mathbb{R}^{d})$ endowed with the Wasserstein distance $W_{2}$, originally exploited in the seminal work by Otto in [Otto], and later developed in the monograph by Ambrosio et al. [AGS], exhibits a singular and non linear structure. As a result, performing differential calculus in $\mathcal{P}_{2}(\mathbb{R}^{d})$ involves the non-trivial task of developing a suitable notion of differentiation in a non-smooth realm. In [AGS], a subdifferential calculus has been introduced, allowing to give a finer description of the Gradient Flows in the space of probability measures. In particular, superdifferentiability of the squared Wasserstein distance is established. We remark that the lack of (local) differentiability of the distance is a common issue in the theory of HJEs in infinite dimensional spaces. The infinite dimensional theory has been firstly studied by Crandall and P.L. Lions in the series of papers [CrandallLions1984, CrandallLions1986a, CrandallLions1986b, CrandallLions1990, CrandallLions1991], where, to treat HJEs in Banach spaces, it is assumed the structural hypothesis of the existence of a smooth metric induced by the assigned norm.
@@
  • Penalization via entropy terms. The singular geometry of the space is reflected by the non-differentiability of the squared Wasserstein distance $W_{2}$ at “non-regular” measures (see [gigc, Section  2]). As described by Daudin and Seeger in [DS], by Bertucci and P.L. Lions in [BertucciApprox], and in the series of papers of the second author et al. [CKTExi],[CKTComp], wellposedness for semilinear HJEs in the space of probability measures can be obtained by adding an entropic penalization that localizes the equation at points of differentiability of the squared distance. The semilinearity assumption is fundamental since it helps to absorb the variational contribution of the entropy penalization in the HJE. These approaches are reminiscent of those introduced by Tataru [Tataru1992, Tataru1994] to treat singular HJBs in Banach spaces, and the one developed by Feng and Katsoulakis [FengKatsoulakis2009] to handle gradient flows in infinite dimensions.
@@
  • Regularization via weaker metrics. The lack of regularity of the distance can be addressed by replacing it with a more regular alternative, following a program analogous to that developed in Banach spaces. One possibility is to modify the geometric structure of the space by introducing weaker yet sufficiently smooth metrics, as discussed by Soner and Yan in [STorus]. A similar idea appears in Cosso et al. [cosso2022masterbellmanequationwasserstein], where a gauge function is employed and common noise is also handled. Along the same lines, Bayraktar et al. [bayraktar2023comparisonviscositysolutionsclass] introduce a clever change of variables that allows one to treat common noise as a finite-dimensional operator, to which an Ishii–Lions lemma (see [crandall1992usersguideviscositysolutions]) can be applied. However, this approach has a drawback: the conditions required to guarantee the comparison principle become difficult to satisfy. Indeed, the weaker the metric, the more regularity is demanded of the data. This trade-off is absent in the penalization approach, where the assumptions on the Hamiltonian closely match those in the finite-dimensional setting.
@@
  • Lifting to a Hilbert space. In the courses at the Collège de France [College], P.L. Lions proposed to lift the HJE to a more familiar and smoother space, such as a Hilbert space, where an extensive viscosity solution theory has been developed in the 80’s and 90’s. Under this viewpoint, the Wasserstein space is realized as a (metric) submersion, and can be regarded as the space of laws of squared-integrable random variables. creates a bridge between the intrinsic calculus developed in [AGS] and extrinsic calculus via the so called Lions’ derivative. See also Carmona and Delarue [CarmonaDelarue, Chapter  5], Bertucci [BSOT], and Gangbo and Tudorascu [GANGBO2019119], where the use of Lions’ lift to treat HJE over $\mathcal{P}_{2}(\mathbb{R}^{d})$ has been deeply treated. For HJEs without idiosyncratic noise terms, this lift provides a natural strategy to give a well-posed notion of viscosity solution in the space of probability measures [BSOT]. However, as already described in [DS, Rmk 2.20], this approach does not adapt well to treat idiosyncratic noise, due to the singularity of the lifted idiosyncratic operator.
@@
  • Viewing the Wasserstein space as a Positively Curved (PC) space. Some authors have instead chosen to develop a more general and abstract theory of HJEs in merely (lenght) metric spaces, see [AMBFeng, GangboSwiech2015_metricViscosity]. Althoug this approach is elegant, it restricts the class of admissible Hamiltonians to the ones depending only on the metric slope, therefore not sensible to directional variations of the unknown function. A further step was to observe that $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a PC space, see [AGS, Appendix]. This geometric property allows one to define a notion of tangent bundle endowed with a weak scalar product. A test-function approach can then be employed to obtain well-posedness. See also[JerhaouiAussedatZidani2024] for an application of this construction. As observed in [DS], the link between test functions and the subdifferential approach is not yet understood. For this reason, in the idiosyncratic case we adopt a more classical notion of solution, leaving the investigation of the relationship between the two frameworks to future work.
@@
  • Superdifferentiability of the Wasserstein distance. In [BSOT], the author realizes that a superdifferentiability of the Wasserstein distance is sufficient to establish a comparison principle for (1.1). In particular, he proposes a notion that ensures both the superdifferentiability of the squared Wasserstein distance and existence results based on the Hilbertian lift.
@@
We describe here the main contributions that we add to the literature:
@@
    Unifying notion of superdifferential:
@@
      The Wasserstein distance squared is superdifferentiable, following the approach in [BSOT].
@@
      $\partial^{+}$ is weaker than the definition of superdifferential given in [BSOT].
@@
      $\partial^{+}$ coincides with the superdifferential adopted in [DS, Definition  2.5] when restricted to measures that are absolutely continuous w.r.t. the Lesbesgue measure.
@@
    where $\partial_{S}$ is the the strong differential, see Definition 2.7, and $\text{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the geometric Tangent space at a measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, see Definition 2.2 Indeed, $\partial^{+}$ is a subset of the superdifferential defined in [BSOT]. On the other hand, in $\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$, the set absolutely continuous measures w.r.t. the Lesbesgue measure, the baricentric projection yields
    $(\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),W_{\mu})\simeq_{\text{Pr}}(\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),\|\cdot\|_{L^{2}_{\mu}})=(\overline{\left\{\nabla\phi\colon\phi\in C^{\infty}_{c}(\mathbb{R}^{d})\right\}}^{L^{2}_{\mu}(\mathbb{R}^{d})},\|\cdot\|_{L^{2}_{\mu}(\mathbb{R}^{d})})\quad\forall\mu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d}).$
    This identification allows to relate our notion of super differential with the one introduced in [DS], namely:
    $\partial^{+}_{S}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\simeq_{\mathrm{Pr}}\partial^{+}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}).$
    We also refer to Lemma 2.12 for more details. Although this result is not new in the literature, see [GANGBO2019119, Theorem  3.14], we provide here an alternative proof using the weak Riemannian formalism introduced in [GigliPhd] and adopted in this work.
    Moreover, in Proposition (2.14) we show that Wasserstein distance squared is superdifferentiable w.r.t to $\partial^{+}$. This property ensures the uniqueness results for the equations (1.1) and (1.2), relying on the previous contributions [BSOT] and [DS].
@@
    $U(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{d})}\left\{\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\right\},$
@@
    and $L:\mathbb{R}^{d}\to\mathbb{R}$ is the Lagrangian associated to the Hamiltonian $H$. Here $\exp_{\mu}:\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to\mathbb{R}$ denotes the exponential map in the Wasserstein space.
    By showing that the above formula provides a representation of the viscosity solution of (1.1) with convex Hamiltonian $H(x,p,\mu)=H(p)$ and a not necessarily differentiable Lagrangian, we generalize the existence result established in [BSOT].
    A related Hopf-Lax representation formula was previously introduced in [GANGBO2019119, Section  5], where first-order time dependent HJEs are considered in the case the Hamiltonian $H$ is convex and satisfies a structural decreasing property, see equation (5.7) there. This condotion ensures a compatibility between their lifted Hamiltonian and the Wasserstein geometry. In that setting, the Hopf-Lax representation, see for instance [GANGBO2019119, Section 5.4], is established in the quadratic case, relying crucially on the aforementioned decreasing property satisfied by the Hamiltonian. By contrast, our result does not require any such structural decreasing assumption of $H$.
@@
    Geodesic Convexity and Idiosyncratic Operator. In Section 4, we inspection the action of the idiosyncratic operator on geodesic convex functions. To this end, we prove a connection between geodesic convexity and the convexity of the flat derivative w.r.t. the $\mathbb{R}^{d}$ variable. Indeed, we prove the following convexity result on mixtures of measures, which, to the best of our knowledge, is new in the literature.
    Lemma 4.6. Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continuous $\Lambda$-geodesically convex function. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then $\forall h\in[0,1]$ the function νF((1-h)μ+ hν) is $h\Lambda$-geodesically convex.
    The main difficulty in proving this result is that, even if $t\mapsto\nu_{t}$ is a $W_{2}$-geodesic, the curve $(1-h)\mu+h\nu_{t}$ is not. Nevertheless, this curve is still induced by transport coupling, so that using [CavSavSod23, Theorem 9.1], where the equivalence between geodesic convexity and total convexity is shown, we obtain a convex like inequality along this curve, which yields the desired conclusion. We remark that the dimensional restriction $d\geq 2$ comes from [CavSavSod23, Theorem 9.1]. Relying on this result, we do not know if our result can be extended to $d=1$.
    As a byproduct of the previous lemma, we obtain the following result
    Corollary 4.10. Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continous $\Lambda$-geodesically convex function. If $\mathcal{F}$ is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ then $x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is convex in $\mathrm{supp}(\mu)$.
    Here we denoted by $\mathcal{D}_{\mu}$ the flat derivative operator at $\mu$. The proof relies on the well known representation formula
    $\mathcal{D}_{\mu}\mathcal{F}(\mu,x):=\lim_{h\to 0}\frac{\mathcal{F}((1-h)\mu+h\delta_{x})-\mathcal{F}(\mu)}{h}\quad(\mu,x)\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathrm{supp}(\mu).$
    The above results are crucial to prove an estimate for the weak action of the idiosyncratic operator introduced in [DS].
    Proposition 4.5. Suppose $d\geq 2$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu^{-1}\in L^{\infty}_{\text{loc}}$ and $\mu\in W^{1,\infty}_{\text{loc}}$. Let $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}$ be a $\Lambda$-geodesically concave functional that is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
@@
    The final result constitutes the main motivation of this article. Namely, showing the convergence of $U^{\varepsilon}$, the viscosity solution of (1.2), with $a(\mu,x)=\text{Id}_{d\times d}$, towards $U$ viscosity solution of (1.1), providing an optimal rate of convergence.
    Theorem 5.1. Under Assumption 2.15, let $U^{\varepsilon}$ and $U$ be bounded viscosity solutions of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}$, and (1.1), respectively. Assume that $U^{\varepsilon}(T,\mu)=U(T,\mu)=\mathcal{G}(\mu)$. Moreover, suppose $U^{\varepsilon}$ to be $W_{1}$-Lipschitz continuous, while $U$ to be $W_{2}$-Lipschitz continuous, Lipschitz in time, and continuous w.r.t. the weak convergence. Then there exists a constant $C\geq 0$ depending only on the data such that

### v2

@@ Abstract.
The aim of this article is twofold. First, we develop a unified framework for viscosity solutions to both first-order Hamilton-Jacobi equations and semilinear Hamilton–Jacobi equations driven by the idiosyncratic operator, defined on the Wasserstein Space. Second, we establish a vanishing-viscosity limit—extending beyond the classical control-theoretic setting—for solutions of semilinear Hamilton–Jacobi equations, proving their convergence to the corresponding first-order solution as the idiosyncratic noise vanishes. Our approach provides an optimal convergence rate.
@@ 1. Introduction
(1.2) $\partial_{t}U(t,\mu)\;+\;\int_{\mathbb{R}^{d}}H\bigl(x,\partial_{\mu}U(t,\mu,x),\mu\bigr)\,d\mu(x)\;-\;\int_{\mathbb{R}^{d}}\mathrm{tr}\!\left[a(\mu,x)\,\nabla\partial_{\mu}U(t,\mu,x)\right]d\mu(x)=0,$
in $0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$, where, $T>0$, $H:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is the Hamiltonian, and $a:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to M_{d\times d}(\mathbb{R})$ is the diffusion matrix that will be assumed to be positive definite. Usually we will consider the final time condition $U(T,\cdot)=\mathcal{G}(\cdot)$ where $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is the terminal datum. Here, $\partial_{\mu}U(t,\mu,x)$ stands for the Wasserstein gradient of $U$ at the point $(t,\mu,x)$.
@@
In this paper, we develop an approach that enables us to treat these two equations using a unified definition of a viscosity solution. In particular, we establish an asymptotic relation between the viscosity solutions of (1.1) and (1.2) through a vanishing-viscosity limit. To the best of our knowledge, this is the first work to provide such a perturbative link without imposing regularity assumptions on the functions $U^{\varepsilon}$, the viscosity solutions to (1.2) when $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$, apart from those inherent to the control setting. In [47], a vanishing viscosity result is obtained also in the presence of a common noise operator, in the convex case; however, that analysis relies on additional regularity assumptions on the solutions.
@@ 1.1. Overview on the existing literature
The study of HJEs on Wasserstein spaces has gained a lot of interest, in these last years, thanks to their connections with Mean Field Control, Mean Field Games, and the associated Master Equation. Mean Field Games theory has been introduced by Lasry and P.L. Lions [41], and independently by Huang, Caines and Malhamé [39]. It has been subsequently developed in Lions’ courses at Collège de France [42] and it has attracted the interest of a large group of researchers. The well-posedness theory for solutions of HJEs in the space of probability measures is one of the objectives of this theory due to its connection with mean-field particle approximation [11, 16], the control of Mckean-Vlasov equation [14, 13], and large deviation theory [34].
In the study of HJEs in $\mathcal{P}_{2}(\mathbb{R}^{d})$, various notions of viscosity solutions have been proposed in the literature (possibly in presence of a common noise) with the shared aim of extending the classical finite-dimensional framework to this more complex setting. However, this new context presents significant challenges. The geometry of $\mathcal{P}_{2}(\mathbb{R}^{d})$ endowed with the Wasserstein distance $W_{2}$, originally exploited in the seminal work by Otto in [43], and later developed in the monograph by Ambrosio et al. [2], exhibits a singular and non linear structure. As a result, performing differential calculus in $\mathcal{P}_{2}(\mathbb{R}^{d})$ involves the non-trivial task of developing a suitable notion of differentiation in a non-smooth realm. In [2], a subdifferential calculus has been introduced, allowing to give a finer description of the Gradient Flows in the space of probability measures. In particular, superdifferentiability of the squared Wasserstein distance is established. We remark that the lack of (local) differentiability of the distance is a common issue in the theory of HJEs in infinite dimensional spaces. The infinite dimensional theory has been firstly studied by Crandall and P.L. Lions in the series of papers [24, 26, 27, 28, 29], where, to treat HJEs in Banach spaces, it is assumed the structural hypothesis of the existence of a smooth metric induced by the assigned norm.
@@
  • Penalization via entropy terms. The singular geometry of the space is reflected by the non-differentiability of the squared Wasserstein distance $W_{2}$ at “non-regular” measures (see [38, Section  2]). As described by Daudin and Seeger in [30], by Bertucci and P.L. Lions in [7], and in the series of papers of the second author et al. [20],[21], wellposedness for semilinear HJEs in the space of probability measures can be obtained by adding an entropic penalization that localizes the equation at points of differentiability of the squared distance. The semilinearity assumption is fundamental since it helps to absorb the variational contribution of the entropy penalization in the HJE. These approaches are reminiscent of those introduced by Tataru [48, 49] to treat singular HJBs in Banach spaces, and the one developed by Feng and Katsoulakis [33] to handle gradient flows in infinite dimensions.
@@
  • Regularization via weaker metrics. The lack of regularity of the distance can be addressed by replacing it with a more regular alternative, following a program analogous to that developed in Banach spaces. One possibility is to modify the geometric structure of the space by introducing weaker yet sufficiently smooth metrics, as discussed by Soner and Yan in [46]. A similar idea appears in Cosso et al. [22], where a gauge function is employed and common noise is also handled. Along the same lines, Bayraktar et al. [6] introduce a clever change of variables that allows one to treat common noise as a finite-dimensional operator, to which an Ishii–Lions lemma (see [25]) can be applied. However, this approach has a drawback: the conditions required to guarantee the comparison principle become difficult to satisfy. Indeed, the weaker the metric, the more regularity is demanded of the data. This trade-off is absent in the penalization approach, where the assumptions on the Hamiltonian closely match those in the finite-dimensional setting.
@@
  • Lifting to a Hilbert space. In the courses at the Collège de France [42], P.L. Lions proposed to lift the HJE to a more familiar and smoother space, such as a Hilbert space, where an extensive viscosity solution theory has been developed in the 80’s and 90’s. Under this viewpoint, the Wasserstein space is realized as a (metric) submersion, and can be regarded as the space of laws of squared-integrable random variables. This lift creates a bridge between the intrinsic calculus developed in [2] and extrinsic calculus via the so called Lions’ derivative. See also Carmona and Delarue [14, Chapter  5], Bertucci [9], and Gangbo and Tudorascu [36], where the use of Lions’ lift to treat HJE over $\mathcal{P}_{2}(\mathbb{R}^{d})$ has been deeply treated. For HJEs without idiosyncratic noise terms, this lift provides a natural strategy to give a well-posed notion of viscosity solution in the space of probability measures [9]. However, as already described in [30, Rmk 2.20], this approach does not adapt well to treat idiosyncratic noise, due to the singularity of the lifted idiosyncratic operator.
@@
  • Viewing the Wasserstein space as a Positively Curved (PC) space. Some authors have instead chosen to develop a more general and abstract theory of HJEs in merely (lenght) metric spaces, see [1, 35]. Although this approach is elegant, it restricts the class of admissible Hamiltonians to the ones depending only on the metric slope, therefore not sensible to directional variations of the unknown function. A further step was to observe that $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a PC space, see [2, Appendix]. This geometric property allows one to define a notion of tangent bundle endowed with a weak scalar product. A test-function approach can then be employed to obtain well-posedness. See also[3] for an application of this construction. As observed in [30], the link between test functions and the subdifferential approach is not yet understood. For this reason, in the idiosyncratic case we adopt a more classical notion of solution, leaving the investigation of the relationship between the two frameworks to future work.
@@
  • Superdifferentiability of the Wasserstein distance. In [9], the author realizes that a superdifferentiability of the Wasserstein distance is sufficient to establish a comparison principle for (1.1). In particular, he proposes a notion that ensures both the superdifferentiability of the squared Wasserstein distance and existence results based on the Hilbertian lift.
@@
We describe here our main contributions, referring to Section 2 for the notations used.
@@
    Unifying notion of superdifferential.
@@
      The Wasserstein distance squared is superdifferentiable, following the approach in [9].
@@
      $\partial^{+}$ is weaker than the definition of superdifferential given in [9].
@@
      $\partial^{+}$ coincides with the superdifferential adopted in [30, Definition  2.5] when restricted to measures that are absolutely continuous w.r.t. the Lesbesgue measure.
@@
    where $\partial_{S}$ is the the strong differential, see Definition 2.7, and $\text{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the geometric Tangent space at a measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, see Definition 2.2 Indeed, $\partial^{+}$ is a subset of the superdifferential defined in [9]. On the other hand, in $\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$, the set absolutely continuous measures w.r.t. the Lesbesgue measure, the baricentric projection Pr, see (2.9), yields $\forall\mu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$
    $(\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),W_{\mu})\simeq_{\text{Pr}}(\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),\|\cdot\|_{L^{2}_{\mu}(\mathbb{R}^{d})})=(\overline{\left\{\nabla\phi\colon\phi\in C^{\infty}_{c}(\mathbb{R}^{d})\right\}}^{L^{2}_{\mu}(\mathbb{R}^{d})},\|\cdot\|_{L^{2}_{\mu}(\mathbb{R}^{d})}).$
    This identification allows to relate our notion of superdifferential with the one introduced in [30], namely:
    $\partial^{+}_{S}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\simeq_{\mathrm{Pr}}\partial^{+}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d}),\quad\forall\mu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d}).$
    We also refer to Lemma 2.12 for more details. Although this result is not new in the literature, see [36, Theorem  3.14], we provide here an alternative proof using the weak Riemannian formalism introduced in [37] and adopted in this work.
    Moreover, in Proposition 2.15 we show that Wasserstein distance squared is superdifferentiable w.r.t to $\partial^{+}$. This property ensures the uniqueness results for the equations (1.1) and (1.2), relying on the previous contributions [9] and [30].
@@
    $U(t,\mu)=\inf_{\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\mathcal{G}(\exp_{\mu}(\gamma))+(T-t)\mathcal{L}\left(\frac{1}{T-t}\cdot\gamma\right)\right\},$
@@
    and $L:\mathbb{R}^{d}\to\mathbb{R}$ is the Lagrangian associated to the Hamiltonian $H$. Moreover, $\exp_{\mu}:\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})\to\mathcal{P}_{2}(\mathbb{R}^{d})$ denotes the exponential map in the Wasserstein space, defined on $\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}):=\left\{\gamma\in\mathcal{P}_{2}{(\mathbb{R}^{2d})},(\pi_{1})_{\#}\gamma=\mu\right\}$.
    By showing that the above formula provides a representation of the viscosity solution of (1.1) with convex Hamiltonian $H(x,p,\mu)=H(p)$ and a not necessarily differentiable Lagrangian, we generalize the existence result established in [9].
    A related Hopf-Lax representation formula was previously introduced in [36, Section  5], where first-order time dependent HJEs are considered in the case the Hamiltonian $H$ is convex and satisfies a structural decreasing property, see equation (5.7) there. This condition ensures compatibility between their lifted Hamiltonian and the Wasserstein geometry. In that setting, the Hopf-Lax representation, see for instance [36, Section 5.4], is established in the quadratic case, relying crucially on the aforementioned decreasing property satisfied by the Hamiltonian. By contrast, our result does not require any such structural decreasing assumption of $H$, see also Remark 3.7.
@@
    Geodesic Convexity and Idiosyncratic Operator. In Section 4, we inspection the action of the idiosyncratic operator on geodesically convex functions. To this end, we prove a connection between geodesic convexity and the convexity of the flat derivative w.r.t. the $\mathbb{R}^{d}$ variable. Indeed, we prove the following convexity result on mixtures of measures, which, to the best of our knowledge, is new in the literature.
    Lemma 4.9. Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continuous $\Lambda$-geodesically convex function. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then $\forall h\in[0,1]$ the function P_2(R^d) νF((1-h)μ+ hν) is $h\Lambda$-geodesically convex.
    The main difficulty in proving this result is that, even if $t\mapsto\nu_{t}$ is a $W_{2}$-constant speed geodesic, the curve $(1-h)\mu+h\nu_{t}$ is not, in general. Nevertheless, this curve is still induced by a transport coupling, so that using [15, Theorem 9.1], where the equivalence between geodesic convexity and total convexity is shown, we obtain a convex like inequality along this curve, which yields the desired conclusion. We remark that the dimensional restriction $d\geq 2$ comes from [15, Theorem 9.1]. Relying on this result, we do not know if our result can be extended to $d=1$.
    As a byproduct of the previous lemma, we obtain the following result.
    Corollary 4.11. Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continuous $\Lambda$-geodesically convex function. If $\mathcal{F}$ is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ then $\mathbb{R}^{d}\ni x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is $\Lambda$-convex.
    In the above, we denoted by $\mathcal{D}_{\mu}$ the flat derivative operator at $\mu$. The proof of the corollary relies on the well known representation formula
    $\mathcal{D}_{\mu}\mathcal{F}(\mu,x):=\lim_{h\to 0}\frac{\mathcal{F}((1-h)\mu+h\delta_{x})-\mathcal{F}(\mu)}{h}\quad(\mu,x)\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}.$
    The above results are crucial to prove an estimate for the weak action of the idiosyncratic operator introduced in [30].
    Proposition 4.7. Suppose $d\geq 2$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ s.t. $\mathcal{I}(\mu)<\infty$, $\mu^{-1}\in L^{\infty}_{\text{loc}}(\mathbb{R}^{d})$ and $\mu\in W^{1,\infty}_{\text{loc}}(\mathbb{R}^{d})$. Let $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}$ be a $\Lambda$-geodesically concave functional that is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then
@@
    The final result constitutes the main motivation of this article. Namely, showing the convergence of $U^{\varepsilon}$, the viscosity solution of (1.2), with $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$, towards $U$ viscosity solution of (1.1), providing an optimal rate of convergence.
    Theorem 5.1. Suppose Assumptions 2.16 hold. Let $U^{\varepsilon}$ and $U$ be two bounded viscosity solutions of (1.2) and (1.1), respectively, s.t. that $U^{\varepsilon}(T,\mu)=U(T,\mu)=\mathcal{G}(\mu)$. Additionally, suppose $U^{\varepsilon}$ to be uniformly in time $W_{1}$-Lipschitz continuous for each $\varepsilon$, and $U$ to be uniformly in time $W_{2}$-Lipschitz continuous. Moreover, suppose $U$ to be jointly continuous with respect to the product topology of $0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$, where $\mathcal{P}_{2}(\mathbb{R}^{d})$ is endowed with the weak convergence as in Definition 2.17. Then, there exists a constant $C\geq 0$ depending only on the data s.t.